### v1

# Enhancing light-matter coupling with electric fields in polar semiconductors

Abstract
Author to whom correspondence should be addressed: Nick Pant, nickpant@umich.edu
*Current Institution: University of Texas at Austin, Austin, TX, USA 78712
Abstract
We revisit the prevailing notion that spontaneous and piezoelectric polarization fields are inherently detrimental to light emission in polar semiconductor heterostructures. Here, we demonstrate that polarization can enhance light-matter coupling in quantum wells through the quantum-confined Stark effect (QCSE), by enabling a reduction in well width without shifting the emission to shorter wavelengths. We illustrate this with the paradigmatic example of red-emitting (644 nm–698 nm) AlInGaN heterostructures, which are essential for achieving polychromatic micron-scale pixels. By systematically exploring their design space with machine learning, we uncover that the globally optimal designs feature extreme internal electric fields exceeding 8 MV/cm. Structural and compositional engineering of electric fields not only offers precise control over the emission wavelength but also boosts the light-emission probability by orders of magnitude.
Significance Statement
Nitride semiconductors are poised to enable micron-scale pixels that underpin novel extended-reality and biomedical technologies. At the same time, this material system is gaining attention for its spontaneous polarization, which can be electrically switched for certain cation compositions. Traditionally, polarization fields have been considered detrimental to light emission, limiting cross-fertilization between these research areas. Our study breaks from this consensus by demonstrating that polarization can substantially enhance light emission. In fact, synthesis limitations in this material system might render electric fields necessary to achieve efficient red-wavelength emission. This discovery conceptually paves the way for novel functionalities that leverage the physics of polarization fields in solid-state lighting and display applications.
preprint: AIP/123-QED
Internal electric fields arise from polarization discontinuities at the interfaces of semiconductors with broken inversion symmetry. Such fields have been used to induce free carriers in wide-bandgap semiconductors without dopant impurities,Chaudhuri et al. (2019) to enable switchable non-volatile memory,Kim et al. (2023) and to separate electrons from holes to improve the performance of solar cells and photocatalysts.Yuan et al. (2011); Frost et al. (2014); Kibria et al. (2016) However, polarization has traditionally been considered detrimental for light-emission applications, such as light-emitting diodes (LED) and lasers, where efficient electron-hole recombination is essential.Fiorentini et al. (9 15); Waltereit et al. (2000); Pimputkar et al. (2009); Dreyer et al. (2016); Auf der Maur et al. (1 15); Lin et al. (2023)
Nevertheless, wurtzite InGaN/GaN quantum wells, featuring strong spontaneous and piezoelectric polarization fields ($>$1 MV/cm), underpin commercial solid-state lighting.Pimputkar et al. (2009) Because the band gap of these LEDs can be tuned across the visible spectrum, they are ideally suited for integrating blue, green, and red wavelengths onto a monolithic platform, crucial for creating miniaturized polychromatic pixels that have applications in extended reality and biomedicine.Lin and Jiang (2020); Behrman and Kymissis (2022) However, improving the energy efficiency of InGaN for red emission remains challenging.Lin et al. (2023)
The drop in efficiency of InGaN LEDs with increasing wavelength is attributed to the field-induced separation of electron and hole wave functions in their quantum wells [Fig. 1a].Auf der Maur et al. (1 15); Lin et al. (2023) Achieving red emission requires high In composition, which increases internal electric fields and reduces radiative recombination, quantified by the $B$ coefficient [Fig. 1b]. Unfortunately, making the quantum wells thinner, to bring electrons and holes closer, shifts the emission to shorter wavelengths, necessitating even higher In compositions to maintain red emission [Fig. 1c]. The challenges associated with achieving high-quality InGaN with over 30% In have been extensively discussed,Schulz et al. (2020); Mishra et al. (2021); Hu et al. (8 23) and is connected to a surface reconstruction during growth that limits In incorporation.Lymperakis et al. (2018) Mitigating the polarization of nitride LEDsWaltereit et al. (2000); Pimputkar et al. (2009); Young et al. (2 08); Lin et al. (2023) has yielded limited success at long wavelength because non-polar designs need more than 40% In for red emission.
Figure 1: Origin of the red efficiency gap in InGaN LEDs. a, Band diagram of InGaN quantum wells, showing how internal polarization fields separate electron and hole wave functions, reducing their overlap and probability of recombination. b, An increase in the In composition increases the polarization field, which dramatically decreases the wave-function overlap. The square of the overlap is correlated with the $B$ coefficient of quantum wells, $B_{\text{QW}}$, relative to their bulk value, $B_{\text{bulk}}$; experimentally measured $B_{\text{QW}}/B_{\text{bulk}}$ in InGaN/GaN single quantum wellsDavid et al. (2019) are shown for qualitative comparison. c, To improve the wave-function overlap within quantum wells, the simplest strategy is to decrease the well width, however this is challenging in conventional InGaN/GaN quantum-wells as thinner wells require In compositions higher than thermodynamically favored, in order to maintain red emission.
Here, we remark that decreasing the quantum-well width can viably enhance light-matter coupling if the emission blueshift can be avoided. In this regard, we propose the enterprising use of electric fields to counterbalance the blueshift through the QCSE, without increasing the In content. Rather than mitigating electric fields, this counter-intuitive strategy exploits the QCSE to enable a reduction in well width, thereby improving the electron-hole overlap.
To evaluate the plausibility of this approach, we systematically explored the configuration space of AlxIn1-x-yGayN quantum wells, restricting the search to realistic compositions of 30% or less In. We addressed the combinatorial challenge of such an exploration through ensemble machine-learning (ML) algorithms that provide a 340$\times$ speedup over quantum-mechanical calculations, reproducing a wide range of experimental measurementsDavid et al. (2019); Xue et al. (2023) without any adjustable parameter.
The red-emitting quantum wells with the strongest light-matter coupling feature extremely large polarization fields, exceeding 8 MV/cm. Their wave-function overlaps are comparable to state-of-the-art blue emitters. We describe how to attain such fields through strategies that have already been demonstrated to be experimentally feasible,Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) namely the incorporation of Al into the barrier and the relaxation of strain within the quantum well. These findings reveal the essential role that electric fields can play in enhancing light-matter coupling.
Results
Machine-learning accelerated search
ML surrogate models have emerged as a promising approach for accelerating expensive quantum-chemical calculations to virtually screen organic moleculesGómez-Bombarelli et al. (2016) and inorganic compounds.Choubisa et al. (2023); Merchant et al. (2023) Here, we use ML to accelerate our exploration of the design space of AlInGaN heterostructures, which is challenging because of the combinatorial explosion of the design space with an increasing number of design variables. We trained our ML models on effective-mass HamiltoniansKohn (1957) derived from first-principles DFT calculations, which we have previously shown can reliably model the spectral and recombination properties of blue- and green-emitting InGaN quantum wells grown at Lumileds.Pant et al. (2022); Li et al. (2023b) See Methods for how we built the Hamiltonian, solved Schrödinger’s and Poisson’s equations, variationally minimized the exciton problem, and tested the speedup afforded by the ML algorithm.
Configuration Emission [eV] Squared overlap
Theory Expt. Theory $B_{\text{QW}}^{\text{expt.}}/B_{\text{bulk}}$
3 nm In0.3Ga0.7N
18 nm GaN
1.90 N/A 0.0069 N/A
3.1 nm In0.29Ga0.71N
1.2 nm Al0.90Ga0.10N
9.6 nm GaN
1.91 1.90 0.0083 0.0051
3.3 nm In0.27Ga0.73N
1.2 nm Al0.82Ga0.18N
9.5 nm GaN
1.97 1.92 0.0056 0.0030
3.5 nm In0.28Ga0.72N
1.25 nm Al0.82Ga0.18N
9.0 nm GaN
1.85 1.92 0.0030 0.0022
3.7 nm In0.28Ga0.72N
1.2 nm Al0.82Ga0.18N
10.4 nm GaN
1.79 1.89 0.0016 0.0011
Table 1: Benchmark of quantum calculations against experiments. Our quantum-mechanical calculations accurately predict the salient features of red-emitting quantum wells reported in the literature,Xue et al. (2023) without any adjustable parameter. Our ML models are trained on data generated by these calculations.
In Table I, we benchmark our quantum-mechanical calculations against experimental measurements of red InGaN emitters from the literature.Xue et al. (2023) Without any free parameter, our calculations predict the experimental photoluminescence energies with a mean absolute error of 0.06 eV, which is within the $\sim$0.1 eV error of the underlying ab initio calculations. Interface roughness at the quantum-well interface could yield $\sim$0.1 eV fluctuations in the emission (discussed in Section 5 of the SI), which could also account for the discrepancy between theory and experiment. Although the squared overlap of the lowest eigenstates is not directly measurable, it correlates with the decrease in the radiative recombination $B$ coefficient of quantum wells, $B_{\text{QW}}$, relative to the bulk value of 6×$\times 10^{-11}$ cm3/s.Kioupakis et al. (2013); Schiavon et al. (2013) The calculated squared overlaps systematically predict the trends in experimental $B_{\text{QW}}/B_{\text{bulk}}$. Additional benchmarks of our calculations against experimentally measured $B$ coefficientsDavid et al. (2019) at different wavelengths are provided in Fig. 1b for InGaN/GaN single quantum wells.
We defined our search space using practical considerations. We considered quantum wells separated from their periodic images by large barriers of 18 nm to, initially, focus on the physics of single quantum wells. We employed $c$-plane GaN substrates reflecting the current standard for LED epitaxy, and considered well widths between 0.3 nm and 4.0 nm, as thicker wells are likely to relax strain by generating defects. Lastly, we required the In composition in both the well and barrier to be 30% or less.
Figure 2: Machine learning for heterostructure exploration. a, Diagram of the ML accelerated workflow: we employed ensemble ML models, trained on first-principles-derived quantum-mechanical data, as surrogates for efficiently screening candidate quantum-well configurations. We filtered the candidates according to the experimental feasibility of their composition, the optical band gap, and the wave-function squared overlap. b, Performance of the ML surrogates for predicting the electronic band gap compared to quantum-mechanical calculations. c, The ML predicted exciton binding energy compared to quantum-mechanical calculations. d, The ML predicted squared overlap of electron and hole wave functions compared to quantum-mechanical calculations.
We illustrate our ML accelerated workflow in Fig. 2a. Initially, a linear-regression model regularized with the least absolute shrinkage and selection operator (LASSO) predicts polarization fields within quantum wells. Subsequently, these predicted fields are cascaded to random decision forests, which predict the target variables by learning non-linear relationships using ensembles of decision trees. We chose these models because they are widely available, computationally inexpensive, and easily trained on outputs of any numerical solver. We provide details in Methods. Fig. 2b,c,d shows predicted vs quantum-mechanically calculated electronic band gap, exciton binding energy, and base-10 logarithm of the wave-function squared overlap. These benchmarks are calculated on a hold-out dataset of 4000 randomly sampled configurations in the entire configuration space, not restricted to In $\leq$ 0.3.
Using the trained surrogate models, we exhaustively searched the configuration space, screening for optical band gaps between 1.775 eV and 1.925 eV, which corresponds to red emission. Initially, we screened for large oscillator strengths by filtering for wave-function squared overlaps that are at least 3$\times$ larger than reference 3 nm wide In0.3Ga0.7N/GaN well. Explicit verification with quantum-mechanical calculations yielded 6591 candidates in total. Out of this pool, 851 candidates feature wave-function squared overlaps at least 5$\times$ larger than the reference; we hereafter refer to these 851 designs as "optimal."
We present a survey of the the optimal quantum wells in Fig. S1 of the Supporting Information. Interestingly, every optimal well is between 1.8 nm and 2.2 nm thick, as thinner wells require In composition greater than 30% to reach 1.85 eV emission and thicker wells are unable to reach 5×5$\times$5 × greater squared wave-function overlap than the reference InGaN/GaN quantum well. Moreover, the wells uniformly contain between 26% to 30% In, with minimal to no Al, while the barriers are rich in Al and sometimes also contain In. Further configurational details can be found in Supporting Dataset 1.
The role of polarization fields
Figure 3: Role of extreme fields in the optimal red-emitting quantum wells. a, Map of the polarization field across an In0.3Ga0.7N well as a function of the AlInGaN barrier composition. The predicted configurations (black points) cluster in a narrow region of configuration space, bounded by contours corresponding to extremely large fields. b, Despite exhibiting significantly larger fields, the predicted quantum wells (black and blue points) exhibit greater wave-function overlap than the reference InGaN/GaN quantum well (red point). “Max field" refers to the In0.3Ga0.7N/AlN configuration, which maximizes the field within our design space. c, Dependence of well width on field strength for In0.3Ga0.7N wells with an optical band gap of 1.82 eV. The black curves correspond to the band gap of polar quantum wells predicted by eq. 1, corresponding to the particle-in-a-triangular-potential model (solid line, $f=1$; dashed line, $f=0.8$; $f$ is an empirical scaling factor that compensates the overestimation of the confinement energies due to the infinite barrier). Extreme fields enable the well width to decrease, improving the optical oscillator strength.
The most striking feature of the optimal configurations is their extremely large polarization fields, ranging from 8 MV/cm to 12 MV/cm. These fields are related to the polarization mismatch between the well and barrier materials. The relationship between barrier composition and polarization fields within the well is best elucidated through a contour map, shown in Fig. 3a, where we approximate the well material as In0.3Ga0.7N since the optimal designs have nearly identical well compositions. Upon superimposing upon this map the barrier compositions of the optimal configurations, it becomes evident that the configurations cluster within a narrow region defined by the contour lines of extremely large polarization fields. We have shown the coordinates of the 2.0 nm-wide optimal quantum wells in Fig. 3a to streamline the illustration, but this analysis applies to other well widths as well.
Despite their extreme fields, the wave-function overlaps in the optimal wells surpass those of the reference well with GaN barriers, as illustrated in Fig. 3b. In Section 2 of the Supporting Information, we show that these conclusions do not change upon accounting for alloy disorder, where we also include the overlap contributions of the excited electron and hole states. Our analysis also suggests that the optimal quantum wells exhibit $B$ coefficients around $\sim 10^{-11}$ cm3/s, comparable to state-of-the-art blue emitters.Li et al. (2023b) Although the Al-rich barriers confine carriers more strongly within the well because of their larger band offsets,Vichi et al. (2020); Xue et al. (2023) it is the polarization fields that cause the primary improvement in wave-function overlap. We discuss our numerical control experiments to probe this hypothesis in Section 3 of the Supporting Information and Fig. S2.
We place this understanding on a general footing by considering, analytically, the band gap change in polar quantum wells due to the QCSE and carrier confinement,Davies (1998)
$\varepsilon_{\text{G}}^{\rm{QW}}=\varepsilon_{\text{G}}^{\rm{bulk}}$ $-eFL_{\text{QW}}$
$+2.338f\Bigg{[}\Big{(}\frac{\hbar^{2}e^{2}F^{2}}{2m_{e}^{*}}\Big{)}^{1/3}+\Big{(}\frac{\hbar^{2}e^{2}F^{2}}{2m_{h}^{*}}\Big{)}^{1/3}\Bigg{]}.$ (1)
Here, $\varepsilon_{\text{G}}^{\rm{QW}}$, represents the quantum-well band gap, while the first term on the left side of the equation, $\varepsilon_{\text{G}}^{\text{bulk}}$, represents the band gap of the bulk material. The second term represents the voltage dropped by the electric field $F$ over a distance $L_{\text{QW}}$, corresponding to the well width. The terms inside the square brackets are confinement energies for electrons ($e$) and holes ($h$) in the triangular potential, bound to one side by the field and the other side by an infinite barrier; the scaling factor $f$ compensates for the overestimation of confinement energies caused by the infinite barrier. With this model, we calculated the field strength and well width pairs that yield quantum-well band gaps of 1.82 eV, if the bulk band gap corresponds to In0.3Ga0.7N. We have neglected the exciton binding energy in this analysis, which is typically one to two orders of magnitude smaller than the band gap. We have plotted the results of these calculations in Fig. 3c, with and without scaling-factor corrections ($f=0.8$ and $f=1$). The role of extreme fields, confirmed by this physically transparent model that connects the reference InGaN/GaN configuration to the optimal configurations, is to enable a reduction in well width without changing the emission wavelength.
As the final verification, we push this concept to its logical extreme by considering In0.3Ga0.7N wells with AlN barriers, a combination that maximizes the internal field within our design space. Detailed calculations reveal a field strength of 13 MV/cm and an optical band gap of 1.82 eV at a well width of 1.62 nm, aligning excellently with the prediction of equation (1) as depicted in Fig. 3c. The wave-function overlap associated with this configuration is indeed an order of magnitude larger than that of 3 nm InGaN/GaN quantum wells, as shown in Fig. 3b. Our initial search space did not encompass well widths of 1.62 nm, reflecting the practical necessity of discretizing continuous variables to numerically represent them. Nevertheless, this limitation did not hinder us from identifying the broader design rule that ultimately led to the discovery of this configuration.
A geometric blueprint for extreme fields
In addition to compositional engineering, there is more freedom in modifying polarization fields by exploring qualitatively different heterostructure designs. We facilitated this exploration by considering InGaN quantum wells separated from GaN barriers with AlInGaN interlayers, a class of heterostructures already employed in state-of-the-art LEDs.Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) We recover the limit of quantum wells with pure AlInGaN barriers, corresponding to our initial search space, by setting the interlayer thickness equal to the barrier thickness. To streamline our understanding, we focused on AlN interlayers, although our conclusions apply to other compositions.
To calculate a geometric map of polarization fields, we solved Poisson’s equation in configurations with 2 nm wide In0.3Ga0.7N wells, and we systematically varied the thickness of the interlayers and the barriers, considering interlayers placed on one or both sides of the wells. We modeled structures that are coherently strained on GaN and also considered partially and fully strain-relaxed wells, reflecting emerging schemes that attempt to improve material quality by relaxing strain.Zhang et al. (2020); Dussaigne et al. (2021); Ji et al. (2023); Pandey et al. (2023) The resulting maps of polarization fields are illustrated in Fig. S3 and S4 of the Supporting Information.
Engineering the heterostructure geometry and relaxing their strain enables the precise control of internal fields. For example, the highest polarization fields correspond to symmetric barriers, which maximize the mismatch in polarization on both sides of the well. Additionally, larger interlayer thicknesses tend to increase the field across the well; increasing the interlayer and barrier thickness simultaneously has the largest impact on amplifying fields. In contrast, making the barrier thicker without also making the interlayer thicker decreases the polarization field. Moreover, we find that engineering strain can also modify the fields. For example, relaxing strain within the well can advantageously enhance the fields because it increases the polarization mismatch between the barrier/interlayer and well regions. Therefore, elastic strain relaxation can be used to redshift the band gap and further decrease the well width to promote wave-function overlap. Examples include the incorporation of In into the substrate to minimize the lattice mismatch with the active region Even et al. (2017) and the electrochemical etching of the substrate to modify its elastic response.Ji et al. (2023) These geometric design rules can be combined with compositional engineering to enhance the wave-function overlaps in quantum wells.
Discussion
Nitride heterostructures are a cornerstone of optoelectronics. In principle, their band gaps can be varied from the deep-ultraviolet (200 nm) to the infrared (2 $\mu$m) through alloying, however historical synthesis limitations have restricted their useful band gaps to green and shorter wavelengths. Exploiting the nitrides’ intrinsic polarization may circumvent this longstanding bottleneck.
Although polarization fields are known to redshift InGaN’s band gap,Vichi et al. (2020) prior studies have fundamentally viewed polarization as a factor that reduces wave-function overlap. To the contrary, by globally mapping the design space of feasible AlInGaN heterostructures, we have demonstrated that designs with the largest wave-function overlaps exhibit extreme fields because they support thinner wells. While it is true that non-polar quantum wells exhibit unity wave-function overlap, they require infeasibly high In concentrations exceeding 40% and well widths substantially thicker than 3 nm for red emission. Promising approaches are emerging that improve In incorporation through strain relaxation,Zhang et al. (2020); Dussaigne et al. (2021); Ji et al. (2023); Pandey et al. (2023) however no planar LED with such high In content and wide wells have ever been reported with associated external quantum efficiency greater than 1%. This suggests that extreme fields may be necessary to achieve efficient red emission.
Notably, our findings align with experimental reports of the most efficient red LEDs, many of which feature Al-containing barriers or strain-relaxed quantum wells.Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) Hence, we suggest critically reevaluating whether polarization fields already play a role in improving the performance of these LEDs. Overall, carriers recombine more quickly in quantum wells with strong light-matter coupling, meaning fewer carriers are needed to achieve high current densities. This helps to reduce the carrier-density-dependent hue shiftPant et al. (2022) and mitigate efficiency droop, a phenomenon where the efficiency decreases with increasing carrier density.Li et al. (2023b)
However, there are potential downsides to increasing the fields. We recently showed that the confinement of carriers by internal fields amplifies intrinsic $C$ coefficients, which quantifies non-radiative Auger-Meitner recombination.Pant et al. (2024) This would necessitate a design trade-off between increasing the $B$ and $C$ coefficients. Increasing the $B$ coefficient would be advantageous, even at the expense of increasing the $C$ coefficient, in materials with high defect density because an increase in the light-emission probability would reduce the fraction of carriers that recombine non-radiatively over defects. An additional complication is that the spectral linewidth may be sensitive to small fluctuations of the well width in the presence of strong fields, which would need to be carefully controlled. We consider the impact of monolayer fluctuations on the emission in Section 5 of the SI.
There are also practical limits to how thick Al-containing layers can be grown on GaN before crack formation. The crack-free thickness for AlGaN on GaN templates was found experimentally to be between 50 and 100 nm for AlN mole fractions between 30% and 40%.Bethoux et al. (2003) Cracking could place constraints on the total number of quantum wells and the maximum barrier thickness in LEDs with Al-rich barriers and thin InGaN wells. However, a significant part of the theoretical design space is experimentally accessible within those constraints. Moreover, we have identified designs that have minimal lattice mismatch; we present nearly-lattice-matched barrier compositions having strain of 1% or less that yield polarization fields of greater than 8 MV/cm in 2 nm In0.3Ga0.7N quantum wells in Supplementary Dataset 2.
We have primarily focused on improving the light emission of quantum wells, however further investigation is needed to integrate these designs into device architectures with high electrical-injection efficiency. One promising approach is the use of V pits, whose side walls have been shown to promote carrier injection across multiple quantum wells, particularly in the presence of Al-containing intermediate layers.Tak et al. (2023) Another advantage of V pits is that they expose semi-polar planes with reduced electric fields, allowing for low turn-on voltages in highly polar LED structures.Li et al. (6 05) The use of V pits in conjunction with extreme-field quantum wells confers the benefits of a semi-polar design for electrical operation while enabling enhanced light-matter coupling through extreme fields. Extreme-field quantum wells are also ideal candidates for alternative architectures such as blue-pump LEDs, where an electrically driven blue LED optically pumps a red-emitting quantum well to achieve polychromatic emission.Damilano and Gil (2015) This offers the possibility of circumventing the historical challenges of electrical injection in conventional LED architectures through entirely novel paradigms. Finally, in Section 6 of the SI, we demonstrate that the proposed quantum wells are less prone to current overflow than conventional designs, despite their thinner wells, because the Al-containing barriers act as electron-blocking layers.
This study challenges the conventional understanding that internal electric fields are inherently detrimental, showing that polarization can be a useful dimension in designing efficient light emitters. Since our findings are grounded in the fundamental principles governing particles in electric fields, our conclusions are broadly applicable to a wide range of polar semiconductor heterostructures. Therefore, these insights can be extended to polar heterostructures in larger chemical spaces, including emerging ferroelectric nitride,Calderon et al. (2023) ferroelectric perovskite,Das et al. (2020) and piezoelectric low-dimensionalSherrell et al. (2022) semiconductors, in which polarization may enable unanticipated optoelectronic functionalities without compromising light-matter coupling.
Materials & Methods
Poisson’s equation
Our calculations of polarization fields are based on the Berry-phase approach to the Modern Theory of Polarization, using improper polarization constantsDreyer et al. (2016) obtained with DFT employing the Heyd-Scuseria-Erznerhof (HSE) hybrid functional.Heyd, Scuseria, and Ernzerhof (2003) The Hartree-Fock mixing parameters are $\alpha=0.31$ for GaN and AlN, and $\alpha=0.25$ for InN, with range separation of 0.2 Å-1, which reproduce the correct band gaps and structural properties. We calculated the strain $\epsilon$ using ab initio lattice and elastic constants,Wright (1997) from which we evaluated the fixed charge using Gauss’s flux theorem, $\rho=-\nabla\cdot(P_{\text{sp}}+e_{31}\epsilon_{11}+e_{31}\epsilon_{22}+e_{33}\epsilon_{33})$, where $P_{\text{sp}}$ is the spontaneous polarization and $e_{ij}$ are piezoelectric coefficients, also calculated with the HSE functional.Dreyer et al. (2016) Polarization fields are then obtained by solving Poisson’s equation.
Schrödinger’s equation
Optical properties of materials can be accurately described by their band-edge electronic structure because carriers populate only a small fraction of the conduction-band bottom and valence-band top. We calculate properties of semiconductor heterostructures by solving the effective-mass Schrödinger equation, $\hat{H}\psi=E\psi$, where $\hat{H}=-\nabla\cdot\left[\frac{\hbar^{2}}{2m^{*}(\textbf{r})}\nabla\right]+V(\textbf{r})+q\phi(\textbf{r})$, is represented with finite differences. Here, $m^{*}$ is the effective mass,Rinke et al. (2008) $\phi$ is the electrostatic potential obtained from the solution of Poisson’s equation, and $V$ is the effective potential. We used band alignments obtained from accurate DFT slab calculations employing the HSE functional,Moses et al. (2011) interpolated to intermediate compositions with composition-dependent bowing coefficients.Moses et al. (2011); Kyrtsos, Matsubara, and Bellotti (2020); Iliopoulos et al. (2008) We accounted for strain effects via first-principles linear deformation potentials calculated with the HSE functional.Yan et al. (2014) We verified our results by calculating wave-function overlaps with full three-dimensional calculations, accounting for alloy disorder and summing over hundreds of eigenstates using nextnano++.Birner et al. (2007) The parameters used in our calculations are provided in Tables S1-S3 of the Supporting Information.
Variational minimization of excitons
We calculated the exciton binding energy by variationally minimizing the exciton Hamiltonian, $\hat{H}_{\text{ex}}=\hat{H}_{e}+\hat{H}_{h}-\frac{\hbar^{2}}{2\mu^{*}}\nabla^{2}_{\bm{\rho}}-\frac{\hbar^{2}}{2(m^{*}_{e}+m^{*}_{h})}\nabla_{\mathbf{R}}^{2}-\frac{e^{2}}{4\pi\epsilon}\frac{1}{\sqrt{\rho^{2}+(z_{e}-z_{h})^{2}}}$, where $\hat{H}_{e}$ and $\hat{H}_{h}$ are the single-particle electron and hole Hamiltonians, $\mu^{*}$ is the reduced effective mass, $\bm{\rho}$ is the in-plane separation vector ($\textbf{r}_{e}^{\perp}-\textbf{r}_{h}^{\perp}$), R is the center of mass coordinate, and $z_{e},z_{h}$ are the out-of-plane coordinates of the electron and hole. Since light emission occurs from excitons with zero center-of-mass momentum, we neglected the fourth term in the exciton Hamiltonian. Excitons in a quantum well with electric fields can be calculated by making the following separable Ansatz for the wave function,Mares and Chuang (1993) $\Psi^{(\lambda)}_{\text{ex}}(\rho,z_{e},z_{h})=\psi_{e}(z_{e})\psi_{h}(z_{h})\phi^{(\lambda)}(\rho),\text{ where }\phi^{(\lambda)}(\rho)=\sqrt{\frac{2}{\pi}}\frac{e^{-\rho/\lambda}}{\lambda}$, $\lambda$ is the variational parameter and $\psi_{e}$ and $\psi_{h}$ are obtained by diagonalizing the single-particle Hamiltonians along the direction of confinement.
Machine-learning prediction
We used ML surrogate models to exhaustively explore the high-dimensional configuration space of nitride quantum wells. We trained our ML models on explicit quantum-mechanical calculations of 5000 randomly sampled quantum-well configurations. We describe our ML workflow in the main text. We employed scikit-learn’s Pedregosa et al. (2011) LASSO regressor and random forests, fitting 100 decision-tree regressors on bootstrapped quantum-well configurations, allowing the nodes to expand until all leaves were pure. We tested the speedup afforded by the ML workflow using the timeit Python library on a laptop with 16GB random-access memory and a 2.6GHz six-core Intel Core i7 processor, with a maximum single-core frequency of 4.50 GHz. Over 70 runs, the ML prediction took 7.28 ms $\pm$ 1.63 ms per run on average, while the quantum-mechanical calculations were 339×339$\times$339 × slower, taking 2.47 s $\pm$ 33.6 ms per run.
Acknowledgements
We thank Kyle Bushick, Mahlet Molla, Emily Oliphant, and Amanda Wang for their feedback. This project was funded by the U.S. Department of Energy, Office of Energy Efficiency and Renewable Energy, under Award No. DE-EE0009163. Computational resources were provided by the National Energy Research Scientific Computing Center, a Department of Energy Office of Science User Facility, supported under Contract No. DEAC0205CH11231. N.P. acknowledges the support of the Natural Sciences & Engineering Research Council of Canada Postgraduate Scholarship.