### v2

# How polarization fields can be beneficial in nitride emitters

Abstract
Author to whom correspondence should be addressed: Nick Pant, nickpant@umich.edu
*Current Institution: University of Texas at Austin, Austin, TX, USA 78712
Abstract
We revisit the prevailing notion that spontaneous and piezoelectric polarization fields are inherently detrimental to light emission in polar semiconductor heterostructures. Here, we demonstrate that polarization can actually enhance the overlap of electron and hole wave functions in III-nitride quantum wells, which are being considered for polychromatic micron-scale pixels in extended-reality and biomedical applications. By systematically exploring the design space of AlInGaN heterostructures with machine learning, we uncover that larger electric fields globally correlate with higher wave-function overlaps for red emitters. The key is the quantum-confined Stark effect, which enables a reduction in well width while avoiding the associated increase in In composition. Our work demonstrates the use of machine learning to reveal unexpected trends for informing the design of semiconductor heterostructures.
preprint: AIP/123-QED
Internal electric fields arise from polarization discontinuities at the interfaces of semiconductors with broken inversion symmetry. Such fields have been used to induce free carriers in wide-bandgap semiconductors without dopant impurities,Chaudhuri et al. (2019) to enable switchable non-volatile memory,Kim et al. (2023) and to separate electrons from holes to improve the performance of solar cells and photocatalysts.Yuan et al. (2011); Frost et al. (2014); Kibria et al. (2016) However, polarization has traditionally been considered detrimental for light-emission applications, such as light-emitting diodes (LED) and lasers, where efficient electron-hole recombination is essential.Fiorentini et al. (9 15); Waltereit et al. (2000); Pimputkar et al. (2009); Dreyer et al. (2016); Auf der Maur et al. (1 15); Lin et al. (2023)
Nevertheless, wurtzite InGaN/GaN quantum wells, featuring strong spontaneous and piezoelectric polarization fields ($>$1 MV/cm), underpin commercial solid-state lighting.Pimputkar et al. (2009) Because the band gap of these LEDs can be tuned across the visible spectrum, they are ideally suited for integrating blue, green, and red wavelengths onto a monolithic platform, crucial for creating miniaturized polychromatic pixels that have applications in extended reality and biomedicine.Lin and Jiang (2020); Behrman and Kymissis (2022) However, improving the energy efficiency of InGaN for red emission remains challenging.Lin et al. (2023)
The drop in efficiency of InGaN LEDs with increasing wavelength is attributed to the field-induced separation of electron and hole wave functions in their quantum wells [Fig. 1a].Auf der Maur et al. (1 15); Lin et al. (2023) Achieving red emission requires high In composition, which increases internal electric fields and reduces radiative recombination, quantified by the $B$ coefficient [Fig. 1b]. Unfortunately, making the quantum wells thinner, to bring electrons and holes closer, shifts the emission to shorter wavelengths, necessitating even higher In compositions to maintain red emission [Fig. 1c].
The challenges associated with achieving high-quality InGaN with over 30% In have been extensively discussed,Schulz et al. (2020); Mishra et al. (2021); Hu et al. (8 23) and is connected to a surface reconstruction during growth that limits In incorporation.Lymperakis et al. (2018) Mitigating the polarization of nitride LEDsWaltereit et al. (2000); Pimputkar et al. (2009); Young et al. (2 08); Lin et al. (2023) has yielded limited success at long wavelength because non-polar designs need more than 40% In for red emission.
Figure 1: Origin of the red efficiency gap in InGaN LEDs. a, Band diagram of InGaN quantum wells, showing how internal polarization fields separate electron and hole wave functions, reducing their overlap and probability of recombination. b, An increase in the In composition increases the polarization field, which dramatically decreases the wave-function overlap. The square of the overlap is correlated with the $B$ coefficient of quantum wells, $B_{\text{QW}}$, relative to their bulk value, $B_{\text{bulk}}$; experimentally measured $B_{\text{QW}}/B_{\text{bulk}}$ in InGaN/GaN single quantum wellsDavid et al. (2019) are shown for qualitative comparison. c, To improve the wave-function overlap within quantum wells, the simplest strategy is to decrease the well width, however this is challenging in conventional InGaN/GaN quantum-wells as thinner wells require In compositions higher than thermodynamically favored, in order to maintain red emission.
Here, we identify red-emitting quantum wells with the highest wave-function overlaps, by systematically exploring the configuration space of AlxIn1-x-yGayN quantum wells, restricting the search to realistic compositions of 30% or less In. We addressed the combinatorial challenge of such an exploration through ensemble machine-learning (ML) algorithms that provide a 340$\times$ speedup over quantum-mechanical calculations, reproducing a wide range of experimental measurementsDavid et al. (2019); Xue et al. (2023) without any adjustable parameter.
The goal of this search is not to focus on the engineering challenge of optimizing LEDs, which depends sensitively on the consideration of specific structural details. Instead, using idealized models, we take a panoramic view and aim to identify overarching design principles that cannot be discovered through laborious experiments limited to exploring a small region of the design space.
To our surprise, increasing the electric field increases the wave-function overlap in red-emitting quantum wells, as the QCSE enables a reduction in well width while keeping the emission wavelength constant. The wave-function overlaps of the identified red emitters are comparable to state-of-the-art blue emitters. We demonstrate that strategies experimentally shown to improve the performance of red emitters, such as incorporating Al into the barrier and relaxing strain within the quantum well, Hwang et al. (6 23); Ozaki, Funato, and Kawakami (2018); Iida et al. (2020); Li et al. (2023a) also enhance electric fields, raising the possibility that state-of-the-art LED designs may already benefit from increased electric fields.
ML surrogate models have emerged as a promising approach for accelerating expensive quantum-chemical calculations to virtually screen organic moleculesGómez-Bombarelli et al. (2016) and inorganic compounds.Choubisa et al. (2023); Merchant et al. (2023) Here, we use ML to accelerate our exploration of the design space of AlInGaN heterostructures, which is challenging because of the combinatorial explosion of the design space with an increasing number of design variables. We trained our ML models on effective-mass HamiltoniansKohn (1957) derived from first-principles DFT calculations, which we have previously shown can reliably model the spectral and recombination properties of blue- and green-emitting InGaN quantum wells grown at Lumileds.Pant et al. (2022); Li et al. (2023b) See Methods in Section 1 of the Supporting Information (SI) for how we built the Hamiltonian, solved Schrödinger’s and Poisson’s equations, variationally minimized the exciton problem, and tested the speedup afforded by the ML algorithm.
Configuration Emission [eV] Squared overlap
Theory Expt. Theory $B_{\text{QW}}^{\text{expt.}}/B_{\text{bulk}}$
3 nm In0.3Ga0.7N
18 nm GaN
1.90 N/A 0.0069 N/A
3.1 nm In0.29Ga0.71N
1.2 nm Al0.90Ga0.10N
9.6 nm GaN
1.91 1.90 0.0083 0.0051
3.3 nm In0.27Ga0.73N
1.2 nm Al0.82Ga0.18N
9.5 nm GaN
1.97 1.92 0.0056 0.0030
3.5 nm In0.28Ga0.72N
1.25 nm Al0.82Ga0.18N
9.0 nm GaN
1.85 1.92 0.0030 0.0022
3.7 nm In0.28Ga0.72N
1.2 nm Al0.82Ga0.18N
10.4 nm GaN
1.79 1.89 0.0016 0.0011
Table 1: Benchmark of quantum calculations against experiments. Our quantum-mechanical calculations accurately predict the salient features of red-emitting quantum wells reported in the literature,Xue et al. (2023) without any adjustable parameter. Our ML models are trained on data generated by these calculations.
As a starting point, we model ideal quantum wells without alloy disorder or well-width fluctuations. By considering these effects, we have verified that they do not change our main conclusions in Sections 3 and 6 of the SI.
In Table I, we benchmark our quantum-mechanical calculations against experimental measurements of red InGaN emitters from the literature.Xue et al. (2023) It is notable that our idealized model predicts the experimental photoluminescence energies with a mean absolute error of 0.06 eV. Investigating the remaining discrepancy in the prediction of the emission energy requires a careful analysis of the experimental structures measured by Xue et al.Xue et al. (2023) While it is important to recognize the limitations of our model in predicting emission energies of specific LEDs, for which microscopic structural details matter, an accuracy within 0.1 eV is nonetheless sufficient to identify global design trends. We were unable to validate our predictions on thinner quantum wells, for which experimental characterization data remain scarce. Thinner quantum wells are expected to feature stronger excitonic effects, which we have accounted for in our emission predictions with a variational approach (see Methods in Section 1 of the SI).
Although the squared overlap of the lowest eigenstates is not directly measurable, we show that it exactly correlates with the decrease in the radiative recombination $B$ coefficient of quantum wells, $B_{\text{QW}}$, relative to the bulk value of 6×$\times 10^{-11}$ cm3/s.Kioupakis et al. (2013); Schiavon et al. (2013) The calculated squared overlaps systematically predict the trends in experimental $B_{\text{QW}}/B_{\text{bulk}}$. Additional benchmarks of our calculations against experimentally measured $B$ coefficientsDavid et al. (2019) at different wavelengths are provided in Fig. 1b for InGaN/GaN single quantum wells.
We defined our search space using practical considerations. We considered quantum wells separated from their periodic images by large barriers of 18 nm to, initially, focus on the physics of single quantum wells. We employed $c$-plane GaN substrates reflecting the current standard for LED epitaxy, and considered well widths between 0.3 nm and 4.0 nm, as thicker wells are likely to relax strain by generating defects. Lastly, we required the In composition in both the well and barrier to be 30% or less.
Figure 2: Machine learning for heterostructure exploration. a, Diagram of the ML accelerated workflow: we employed ensemble ML models, trained on first-principles-derived quantum-mechanical data, as surrogates for efficiently screening candidate quantum-well configurations. We filtered the candidates according to the experimental feasibility of their composition, the optical band gap, and the wave-function squared overlap. b, Performance of the ML surrogates for predicting the electronic band gap compared to quantum-mechanical calculations. c, The ML predicted exciton binding energy compared to quantum-mechanical calculations. d, The ML predicted squared overlap of electron and hole wave functions compared to quantum-mechanical calculations.
We illustrate our ML accelerated workflow in Fig. 2a. Initially, a linear-regression model regularized with the least absolute shrinkage and selection operator (LASSO) predicts polarization fields within quantum wells. Subsequently, these predicted fields are cascaded to random decision forests, which predict the target variables by learning non-linear relationships using ensembles of decision trees. We chose these models because they are widely available, computationally inexpensive, and easily trained on outputs of any numerical solver. We provide details in Methods. Fig. 2b,c,d shows predicted vs quantum-mechanically calculated electronic band gap, exciton binding energy, and base-10 logarithm of the wave-function squared overlap. These benchmarks are calculated on a hold-out dataset of 4000 randomly sampled configurations in the entire configuration space, not restricted to In $\leq$ 0.3.
Figure 3: Role of extreme fields in the optimal red-emitting quantum wells. a, Map of the polarization field across an In0.3Ga0.7N well as a function of the AlInGaN barrier composition. The predicted configurations (black points) cluster in a narrow region of configuration space, bounded by contours corresponding to extremely large fields. b, Despite exhibiting significantly larger fields, the predicted quantum wells (black and blue points) exhibit greater wave-function overlap than the reference InGaN/GaN quantum well (red point). “Max field" refers to the In0.3Ga0.7N/AlN configuration, which maximizes the field within our design space. c, Dependence of well width on field strength for In0.3Ga0.7N wells with an optical band gap of 1.82 eV. The dashed curve is a guide to the eye. Increasing the electric field enables the well width to decrease, improving the optical oscillator strength.
Using the trained surrogate models, we exhaustively searched the configuration space, screening for optical band gaps between 1.775 eV (698 nm) and 1.925 eV (644 nm), which corresponds to red emission. Initially, we screened for large oscillator strengths by filtering for wave-function squared overlaps that are at least 3$\times$ larger than reference 3 nm wide In0.3Ga0.7N/GaN well. Explicit verification with quantum-mechanical calculations yielded 6591 candidates in total. Out of this pool, 851 candidates feature wave-function squared overlaps at least 5$\times$ larger than the reference; we hereafter refer to these 851 designs as "optimal."
We present a survey of the the optimal quantum wells in Fig. S1. Interestingly, every optimal well is between 1.8 nm and 2.2 nm thick, as thinner wells require In composition greater than 30% to reach 1.85 eV emission and thicker wells are unable to reach 5×5$\times$5 × greater squared wave-function overlap than the reference InGaN/GaN quantum well. Moreover, the wells uniformly contain between 26% to 30% In, with minimal to no Al, while the barriers are rich in Al and sometimes also contain In. Further configurational details can be found in Supporting Dataset 1.
The most striking feature of the optimal configurations is their extremely large polarization fields, ranging from 8 MV/cm to 12 MV/cm. These fields are related to the polarization mismatch between the well and barrier materials. The relationship between barrier composition and polarization fields within the well is best elucidated through a contour map, shown in Fig. 3a, where we approximate the well material as In0.3Ga0.7N since the optimal designs have nearly identical well compositions. Upon superimposing upon this map the barrier compositions of the optimal configurations, it becomes evident that the configurations cluster within a narrow region defined by the contour lines of extremely large polarization fields. We have shown the coordinates of the 2.0 nm-wide optimal quantum wells in Fig. 3a to streamline the illustration, but this analysis applies to other well widths as well.
Despite their extreme fields, the wave-function overlaps in the optimal wells surpass those of the reference well with GaN barriers, as illustrated in Fig. 3b. Our analysis also suggests that the optimal quantum wells exhibit $B$ coefficients around $\sim 10^{-11}$ cm3/s, comparable to state-of-the-art blue emitters.Li et al. (2023b) Although the Al-rich barriers confine carriers more strongly within the well because of their larger band offsets,Vichi et al. (2020); Xue et al. (2023) it is the polarization field rather than improved carrier confinement that causes the primary improvement in wave-function overlap. We discuss our numerical control experiments to probe this hypothesis in Section 4 of the SI and Fig. S2.
As the final verification, we push this concept to its logical extreme by considering In0.3Ga0.7N wells with AlN barriers, a combination that maximizes the internal field within our design space. Detailed calculations reveal a field strength of 13 MV/cm and an optical band gap of 1.82 eV at a well width of 1.62 nm. The wave-function overlap associated with this configuration is indeed an order of magnitude larger than that of 3 nm InGaN/GaN quantum wells, as shown in Fig. 3b. Our initial search space did not encompass well widths of 1.62 nm, reflecting the practical necessity of discretizing continuous variables to numerically represent them. Nevertheless, this limitation did not hinder us from identifying the broader design rule that ultimately led to the discovery of this configuration.
In addition to compositional engineering, there is more freedom in modifying polarization fields by exploring qualitatively different heterostructure designs. We facilitated this exploration by considering InGaN quantum wells separated from GaN barriers with AlInGaN interlayers, a class of heterostructures already employed in state-of-the-art LEDs.Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) We recover the limit of quantum wells with pure AlInGaN barriers, corresponding to our initial search space, by setting the interlayer thickness equal to the barrier thickness. To streamline our understanding, we focused on AlN interlayers, although our conclusions apply to other compositions.
To calculate a geometric map of polarization fields, we solved Poisson’s equation in configurations with 2 nm wide In0.3Ga0.7N wells, and we systematically varied the thickness of the interlayers and the barriers, considering interlayers placed on one or both sides of the wells. We modeled structures that are coherently strained on GaN and also considered partially and fully strain-relaxed wells, reflecting emerging schemes that attempt to improve material quality by relaxing strain.Zhang et al. (2020); Dussaigne et al. (2021); Ji et al. (2023); Pandey et al. (2023) The resulting maps of polarization fields are illustrated in Fig. S3 and S4.
Engineering the heterostructure geometry and relaxing their strain enables the precise control of internal fields. For example, the highest polarization fields correspond to symmetric barriers, which maximize the mismatch in polarization on both sides of the well. Additionally, larger interlayer thicknesses tend to increase the field across the well; increasing the interlayer and barrier thickness simultaneously has the largest impact on amplifying fields. In contrast, making the barrier thicker without also making the interlayer thicker decreases the polarization field.
Moreover, we find that engineering strain can also modify the fields. For example, relaxing strain within the well can advantageously enhance the fields because it increases the polarization mismatch between the barrier/interlayer and well regions. Therefore, elastic strain relaxation can be used to redshift the band gap and further decrease the well width to promote wave-function overlap. Examples include the incorporation of In into the substrate to minimize the lattice mismatch with the active region Even et al. (2017) and the electrochemical etching of the substrate to modify its elastic response.Ji et al. (2023) These geometric design rules can be combined with compositional engineering to enhance the wave-function overlaps in quantum wells.
Polarization fields have been known to redshift InGaN’s band gap and prior works have attempted to leverage this effect to improve the performance of LEDs.Vichi et al. (2020); Koleske et al. (2015) However, this has received limited attention because polarization has been fundamentally viewed as a factor that reduces wave-function overlap. To the contrary, by globally mapping the design space of feasible AlInGaN heterostructures, we have demonstrated that designs with the largest wave-function overlaps exhibit large electric fields because they support thinner wells.
While it is true that non-polar quantum wells exhibit unity wave-function overlap, they require infeasibly high In concentrations exceeding 40% and well widths substantially thicker than 3 nm for red emission. Promising approaches are emerging that improve In incorporation through strain relaxation,Zhang et al. (2020); Dussaigne et al. (2021); Ji et al. (2023); Pandey et al. (2023) however no planar LED with such high In content and wide wells have ever been reported with associated external quantum efficiency greater than 1%.
Notably, our findings align with experimental reports of the most efficient red LEDs, many of which feature Al-containing barriers or strain-relaxed quantum wells.Hwang et al. (6 23); Iida et al. (2020); Li et al. (2023a) Hence, we suggest critically reevaluating whether polarization fields already play a role in improving the performance of these LEDs. Overall, carriers recombine more quickly in quantum wells with strong light-matter coupling, meaning fewer carriers are needed to achieve high current densities. This helps to reduce the carrier-density-dependent hue shiftPant et al. (2022) and mitigate efficiency droop, a phenomenon where the efficiency decreases with increasing carrier density.Li et al. (2023b)
The extent to which polarization can be further useful in practically designing efficient light emitters remains an open question. We have not explicitly modeled complex effects such as well-width fluctuations, inter-well carrier transport, and device-level characteristics, the predictive modeling of which remains a formidable and ongoing challenge. Moreover, there are significant engineering challenges that large electric fields present in LEDs; we discuss these and possible solutions in Section 7 of the SI.
Nevertheless, we have provided clear evidence that internal fields do not need to be mitigated to enhance wave-function overlap. This is compelling given that polarization fields are the reason why thin InGaN quantum wells ($\leq 3$ nm) with low In concentrations ($\leq$ 30%), used in red LEDs, can emit red light in the first place.
III-nitrides are a cornerstone of optoelectronics. With their distinct crystal structure in the III-V family and their proliferation in industry, they offer an unmatched platform to harness polarization effects. While we do not claim to solve the non-trivial challenges associated with further amplifying electric fields, we believe that the challenges are equally great, if not greater, for designs that mitigate these fields. This is a salient point to consider as the community considers moving towards nonpolar designs.Lin et al. (2023) The ability to exploit the unique polarization degree of freedom in the design of optoelectronics should not be underestimated.
Acknowledgements
We thank Kyle Bushick, Mahlet Molla, Emily Oliphant, and Amanda Wang for their feedback. This project was funded by the U.S. Department of Energy, Office of Energy Efficiency and Renewable Energy, under Award No. DE-EE0009163. Computational resources were provided by the National Energy Research Scientific Computing Center, a Department of Energy Office of Science User Facility, supported under Contract No. DEAC0205CH11231. N.P. acknowledges the support of the Natural Sciences & Engineering Research Council of Canada Postgraduate Scholarship.